Every 4 hours, a factory produces 27 cars. How many cars does the factory produce in 11 hours?
Answer by amarjeeth123(574)

Every 4 hours, a factory produces 27 cars.
In 11 hours the factory produces 11*27/4=297/4=74.25 cars
Answer=74.25 cars i.e. approximately 74 cars.